OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. If a = x and x = b, then a = b

OpenStudy (anonymous):

Transitive Property. Here are some of the most used properties. Commutative: a + b + c = b + a + c Associative: a + (b + c) = (a + b) + c Reflexive: If a = b, then a = b Symmetric: If a = b, then b = a Transitive: If a = b, and b = c, then a = c